# RADAR—Radio Afterglow Detection and AI-driven Response: A Federated Framework for Gravitational Wave Event Follow-Up

### v1

@@ 1 Introduction
Here, we present RADARRadio Afterglow Detection and AI-driven Response, a framework for radio follow-up of GW events that enables collaborative, model-informed observing strategies while accommodating heterogeneous data-sharing policies. Proprietary periods vary significantly across facilities, ranging from immediate public release (e.g., some space-based missions) to restricted-access periods in the case of, e.g., LIGO and U.S.  national radio observatories. Data sharing practices also vary among user groups, with some teams releasing measurements promptly without accompanying data, and others delaying dissemination until peer-reviewed publication. Leveraging a privacy-enhancing federated framework, we perform federated learning on proprietary, site-local high-level observables (e.g., flux densities) without transferring raw data. Although RADAR does not implement privacy-preserving federated learning (Li et al., 2025), it is guided by its underlying principles. Rather than training shared AI models (Li et al., 2024), we coordinate light curve fitting through global parameter exchange, facilitating collaborative afterglow characterization while preserving data locality, safeguarding privacy, and respecting institutional data ownership. In this sense, RADAR adopts a privacy-enhancing framework, and enables coordinated, data-informed follow-up campaigns while preserving each group’s autonomy to publish complete analyses, which typically contain data reduction procedures, imaging, localization, spectral energy distribution analysis, and more comprehensive modeling.
As a case study, we apply RADAR to GW170817. Although in that case some key observables (e.g., counterpart position and flux density at discovery) were made promptly available to the community, we show that a federated, privacy-enhancing framework that also incorporates proprietary observations could have enabled a more automated and coordinated effort. Overall, the approach proposed here facilitates broad participation in radio follow-up efforts, even under diverse data-sharing practices, and is readily extendable to other messengers and EM bands.
@@ 2 Computational framework
    The GW module, represented by the left box in Figure 1, is composed of an updated version of the AI models for signal detection previously presented in Tian et al. (2024). Specifically, AI models are sent to the emulated data sources, where LIGO Hanford and Livingston data streams are hosted on the DeltaAI supercomputer at NCSA and the Polaris supercomputer at ALCF, respectively, and upon processing GW data, embeddings (i.e., vector representations of information) are sent to a central server on the Delta supercomputer at NCSA where post-processing is done in search for GW signals. If a GW signal is identified, its GPS time is recorded, and the second component of the analysis is triggered (see below).
@@ 5 The MMA module
The MMA module (bottom box in Figure 1) is designed to derive combined constraints on parameters of the source (binary neutron star merger) using the GW and EM (radio) data. In this work, we emphasize the ability of radio observations to reduce the degeneracy between the luminosity distance $D_{L}$ and the observing angle $\theta_{\rm obs}$ (i.e., the polar angle between the observer’s line of sight and the perpendicular to the plane containing the merging neutron star system) that affects GW observations. Reducing this degeneracy has important consequences, such as, for our ability to use MMA observations to derive standard siren constraints on the Hubble constant $H_{0}$.
@@
The MMA module also performs an overlap analysis by comparing, for the parameters that can be constrained by both the GW and EM data, the posterior distributions from Dingo-BNS and that are derived as output of the radio module. Example results of the Dingo-BNS analysis for GW170817 are presented in Figure 3 and Figure 4. We describe in more detail how the GW and radio posteriors are aggregated by the MMA module in Section 6. In Section 7, we describe the results of this aggregation for the case of GW170817.
@@ 9 Summary and conclusion
We have presented RADAR, a framework aimed at addressing key challenges in the EM follow-up of GW events. Critical among these challenges are the large GW localization regions, which contrast sharply with the comparatively narrow fields of view of sensitive radio arrays, and the requirement for prolonged radio monitoring extending years post-merger to derive accurate astrophysical constraints on the emitting sources and their properties. The demanding observing resource requirements imposed by these two challenges can be mitigated through enhanced global coordination among observers and telescopes. RADAR seeks to advance this objective by enabling community-driven information sharing, federated data analysis, and system resilience, while incorporating AI techniques for both GW signal identification and radio data aggregation and analysis.
@@
Using GW170817 as a case study, we have shown that it is possible to maintain data ownership and privacy while enabling the sharing of predictive models that can guide and optimize follow-up strategies. Although our focus has been on joint GW and radio analyses, the framework presented here is flexible and could be readily extended to a variety of triggers identified in other bands of the EM spectrum (e.g., Möller et al., 2021; Sedgewick et al., 2025; Becker et al., 2025), as well as events in other bands of the GW spectrum (e.g., Yuan et al., 2021; Mangiagli et al., 2022; Amaro-Seoane et al., 2023; Xin & Haiman, 2024; Porter et al., 2025, and references therein). It can also be adapted to incorporate diverse model predictions tailored to different users and scientific objectives. While we have begun to explore the resilience of the framework, a more comprehensive evaluation of potential failure points will be essential to ensure its broader applicability.
An important evaluation metric for AI-driven multi-messenger frameworks such as RADAR is the latency with which data from heterogeneous observatories are acquired, processed, and fused to yield scientifically actionable outputs. This metric is critical for enabling rapid follow-up observations, including the tantalizing possibility of pre-merger alerts that could capture prompt emission from binary neutron star mergers—phenomena that remain largely unexplored (e.g., Chu et al., 2016, and references therein).
In the GW170817 case study, RADAR demonstrated an end-to-end latency on the order of a few hours. However, this aggregate timescale masks the varying latencies of individual pipeline components, each operating within expected bounds. The GW module exemplifies low-latency performance: it processed 4096 s of data from LIGO Hanford and Livingston, sampled at 4096 Hz, in under 4.5 minutes using a batch size of 8 s. The GCN parser completed its task in about 4 minutes, while the Dingo-BNS module required less than 10 minutes to download relevant data and infer $(\theta_{\rm obs},D_{L})$. Communication across modules via Octopus typically took less than 2–3 seconds.
@@
By contrast, the dominant contributor to overall latency was the modeling of the radio data, which involves federated, multi-site fitting of slowly evolving afterglow light curves. Using a distributed log-likelihood MCMC approach, each of 32 walkers proposes a $(\theta_{\rm obs},D_{L})$ pair. These are dispatched to eight geographically distributed radio data sites via Octopus. Each site evaluates its local log-likelihood given the proposed parameters, and the collective results are aggregated to compute the global posterior. This process, which takes approximately 3–4 hours, reflects the inherent timescale of analyzing radio transients that evolve over days to years. Notably, even in a non-federated, centralized setting, this latency would persist due to the cadence and nature of radio observations, which often require $\sim$1 hour per integration.
@@ Appendix B AI-powered GCN parser
Table 4: Summary of the observations related to GW170817 as reported in the GCN circulars. Column 1 lists the GCN circular number. Column 2 provides either the observation time (when explicitly stated) or the GCN publication timestamp. Column 3 records the name of the candidate electromagnetic (EM) counterpart, when available. Column 4 indicates the observing frequency. Column 5 reports the measured flux density or, in the absence of a detection, the upper limit on the flux density. Columns 6 and 7 give the sky coordinates (Right Ascension and Declination in hh:mm:ss and dd:mm:ss, respectively) of the candidate EM counterpart, when provided. Column 8 shows the number of flags (see Section 7 for a detailed description); ‘NA’ marks entries for which GPT 4.1 never returned the ground truth even over multiple runs. A superscript asterisk () denotes a radio detection; all other entries represent non-detections (upper limits). The values in Columns 1–7 represent what would be produced by a trained expert through manual extraction and curation of GW170817-related GCN circulars.

### v2

@@ 1 Introduction
Here, we present RADARRadio Afterglow Detection and AI-driven Response, a framework for radio follow-up of GW events that enables collaborative, model-informed observing strategies while accommodating heterogeneous data-sharing policies. RADAR is not intended to replace current or future low-latency GW pipelines developed and deployed within the LIGO-Virgo-KAGRA collaboration (e.g., Chaudhary et al., 2024, and references therein). Rather, it is a framework complimentary to LIGO-Virgo-KAGRA and other community efforts (e.g., Shawhan, 2012; LIGO Scientific Collaboration et al., 2012; Abbott et al., 2018, 2019; Chatterjee et al., 2019; Gebhard et al., 2019; Sachdev et al., 2020; Magee et al., 2021; Mishra et al., 2021; Williams et al., 2021; Gabbard et al., 2022; Hu & Veitch, 2023; Berbel et al., 2024; Chatterjee et al., 2024; Ewing et al., 2024; Raza et al., 2024; Vallero et al., 2024; Cuoco et al., 2025; Dax et al., 2025; Marx et al., 2025). Instead, RADAR explores how site-local AI inference and federated coordination mechanisms can support cross-observatory workflows under real-world constraints, particularly in the radio domain. The motivation for RADAR stems from the recognition that while centralized alert systems and optical brokers have made significant progress in coordinating follow-up (e.g., Patterson et al., 2019; Barthelmy et al., 2000, 2022; Nordin et al., 2019; Förster et al., 2021; Matheson et al., 2021; Barthelmy et al., 2023; Coughlin et al., 2023; Hosseinzadeh et al., 2024; Coulter et al., 2023; Humensky et al., 2024), the radio community faces distinct challenges that we have already underlined: from narrow fields of view, to extended follow-up needs where rapid contextualization in the form of early radio data modeling may help guide whether long-term monitoring is warranted in resource-constrained settings. RADAR addresses this by enabling model-informed, asynchronous follow-up planning across distributed facilities, without requiring centralized data transfer or early public release. In fact, proprietary periods vary significantly across facilities, ranging from immediate public release (e.g., some space-based missions) to restricted-access periods in the case of, e.g., LIGO and U.S.  national radio observatories. Data sharing practices also vary among user groups, with some teams releasing measurements promptly without accompanying data, and others delaying dissemination until peer-reviewed publication. Leveraging a privacy-enhancing federated framework, we perform federated learning on proprietary, site-local high-level observables (e.g., flux densities) without transferring raw data. Although RADAR does not implement privacy-preserving federated learning (Li et al., 2025), it is guided by its underlying principles. Rather than training shared AI models (Li et al., 2024), we coordinate light curve fitting through global parameter exchange, facilitating collaborative afterglow characterization while preserving data locality, safeguarding privacy, and respecting institutional data ownership. In this sense, RADAR adopts a privacy-enhancing framework, and enables coordinated, data-informed follow-up campaigns while preserving each group’s autonomy to publish complete analyses, which typically contain data reduction procedures, imaging, localization, spectral energy distribution analysis, and more comprehensive modeling.
As a case study, we apply RADAR to GW170817. At the time of writing, GW170817 remains the only binary neutron star merger detected in GWs with a secure EM counterpart identification. GW170817 enjoys a rich MMA dataset which allows controlled, reproducible testing of RADAR across diverse data contributors. Hence, this event serves as a proof-of-concept validation of RADAR’s architecture, including its AI modules, federated orchestration, and MMA integration. Although in the case of GW170817 some key observables (e.g., counterpart position and flux density at discovery) were made promptly available to the community, we show that a federated, privacy-enhancing framework that also incorporates proprietary observations could have enabled a more automated and coordinated effort. Overall, the approach proposed here facilitates broad participation in radio follow-up efforts, even under diverse data-sharing practices, and is readily extendable to other messengers and EM bands. Indeed, RADAR framework’s design allows its modules, whether for GW signal detection, parameter estimation, or radio light curve modeling, to be easily substituted or extended, enabling integration with state-of-the-art developments in each subdomain. Hence, RADAR’s major contribution lies not in proposing new detection models per se, but in integrating key components such as detection, inference, and follow up within a privacy-aware, federated workflow that respects data ownership and adapts to heterogeneous infrastructure. By demonstrating that these components can work together in practice across supercomputing centers, private observatories, and public alert systems, RADAR fills a gap in the current MMA ecosystem: a lightweight, extensible architecture for collaborative science under evolving data-sharing norms.
@@ 2 Computational framework
    The GW module, represented by the left box in Figure 1, is composed of an updated version of the AI models for signal detection previously presented in Tian et al. (2024). Specifically, AI models are sent to the emulated data sources, where LIGO Hanford and Livingston data streams are hosted on the DeltaAI supercomputer at NCSA and the Polaris supercomputer at ALCF, respectively, and upon processing GW data, embeddings (i.e., vector representations of information) are sent to a central server on the Delta supercomputer at NCSA where post-processing is done in search for GW signals. We note that this particular choice of GW detection algorithm is intended solely for demonstration purposes; alternative GW detection pipelines could be integrated within the RADAR framework, as discussed in Section 1. If a GW signal is identified, its GPS time is recorded, compared with that of any public LIGO-Virgo-KAGRA alerts, and the second component of the analysis is triggered (see below).
@@ 5 The MMA module
The MMA module (bottom box in Figure 1) is designed to derive combined constraints on parameters of the source (binary neutron star merger) using the GW and EM (radio) data. In this work, we emphasize the ability of radio observations to reduce the degeneracy between the luminosity distance $D_{L}$ and the observing angle $\theta_{\rm obs}$ (i.e., the polar angle between the observer’s line of sight and the perpendicular to the plane containing the merging neutron star system) that affects GW observations. As already demonstrated by (Hotokezaka et al., 2019) (see also references therein), reducing this degeneracy has important consequences, such as, for our ability to use MMA observations to derive standard siren constraints on the Hubble constant $H_{0}$.
@@
Then, the module performs an overlap analysis by comparing, for the parameters that can be constrained by both the GW and EM data, the posterior distributions from Dingo-BNS and that are derived as output of the radio module. Example results of the Dingo-BNS analysis for GW170817 are presented in Figure 3 and Figure 4. We describe in more detail how the GW and radio posteriors are aggregated by the MMA module in Section 6. In Section 7, we describe the results of this aggregation for the case of GW170817.
We stress that Bayesian inference based on the combination of GW and EM data is not new, and the $H_{0}$-$D_{L}$ inference case presented here is just one possible example of its applications. However, RADAR’s novelty lies the federated orchestration of Bayesian inference, allowing model-driven analysis across private/public, geographically distributed datasets without raw data sharing.
@@ 9 Summary and conclusion
We have presented RADAR, a framework aimed at addressing key challenges in the EM follow-up of GW events. Critical among these challenges are the large GW localization regions, which contrast sharply with the comparatively narrow fields of view of sensitive radio arrays, and the requirement for prolonged radio monitoring extending years post-merger to derive accurate astrophysical constraints on the emitting sources and their properties. The demanding observing resource requirements imposed by these two challenges can be mitigated through enhanced global coordination among observers and telescopes. RADAR seeks to advance this objective by enabling community-driven information sharing, federated data analysis, and system resilience, while incorporating AI techniques for both GW signal identification and radio data aggregation and analysis. We note that RADAR can also be employed to facilitate follow up observations and modeling efforts within multiple networks of collaborations, not necessarily as a single centralized effort.
@@
Using GW170817 as a case study, we have shown that it is possible to maintain data ownership and privacy while enabling the sharing of predictive models that can guide and optimize follow-up strategies. Although our focus has been on joint GW and radio analyses, the framework presented here is flexible and could be readily extended to a variety of triggers identified in other bands of the EM spectrum (e.g., Möller et al., 2021; Sedgewick et al., 2025; Becker et al., 2025), as well as events in other bands of the GW spectrum (e.g., Yuan et al., 2021; Mangiagli et al., 2022; Amaro-Seoane et al., 2023; Xin & Haiman, 2024; Porter et al., 2025, and references therein). It can also be adapted to incorporate diverse model predictions tailored to different users and scientific objectives. While we have begun to explore the resilience of the framework, a more comprehensive evaluation of potential failure points will be essential to ensure its broader applicability. We envision this to become possible during future observing runs of ground-based GW detectors, such as the LIGO-Virgo-KAGRA O5 (Abbott et al., 2018), as binary neutron star mergers localized to $\lesssim 100$ deg2 should come into reach and optical facilities such as Rubin could enable the identification (and more accurate localization) of their associated kilonovae (Andreoni et al., 2022). This will open the way to a hunt for radio afterglows similar to that of GW170817, something the community has been waiting for since the O2 run.
@@
In the longer-term future, when improved GW detectors may issue pre-merger alerts providing GW localization areas bettern than 100 deg2 at least 5 min before merger (see e.g., Figure in 6 Corsi et al., 2024b), an important evaluation metric for AI-driven multi-messenger frameworks such as RADAR will become the latency with which data from heterogeneous observatories are acquired, processed, and fused to yield scientifically actionable outputs. This metric is critical for enabling rapid follow-up observations, including the tantalizing possibility of pre-merger alerts that could capture prompt emission from binary neutron star mergers—phenomena that remain largely unexplored (e.g., Chu et al., 2016, and references therein). In the GW170817 case study, RADAR demonstrated an end-to-end latency on the order of a few hours. However, this aggregate timescale masks the varying latencies of individual pipeline components, each operating within expected bounds. The GW module exemplifies low-latency performance: it processed 4096 s of data from LIGO Hanford and Livingston, sampled at 4096 Hz, in under 4.5 minutes using a batch size of 8 s. The GCN parser completed its task in about 4 minutes, while the Dingo-BNS module required less than 10 minutes to download relevant data and infer $(\theta_{\rm obs},D_{L})$. Communication across modules via Octopus typically took less than 2–3 seconds. By contrast, the dominant contributor to overall latency was the modeling of the radio data, which involves federated, multi-site fitting of slowly evolving afterglow light curves. Using a distributed log-likelihood MCMC approach, each of 32 walkers proposes a $(\theta_{\rm obs},D_{L})$ pair. These are dispatched to eight geographically distributed radio data sites via Octopus. Each site evaluates its local log-likelihood given the proposed parameters, and the collective results are aggregated to compute the global posterior. This process, which takes approximately 3–4 hours, reflects the inherent timescale of analyzing radio transients that evolve over days to years. Notably, even in a non-federated, centralized setting, this latency would persist due to the cadence and nature of radio observations, which often require $\sim$1 hour per integration.
@@ Appendix B AI-powered GCN parser
Table 4: Summary of the observations related to GW170817 as reported in the GCN circulars. Column 1 lists the GCN circular number. Column 2 provides either the observation time (when explicitly stated) or the GCN publication timestamp. Column 3 records the name of the candidate electromagnetic (EM) counterpart, when available. Column 4 indicates the observing frequency. Column 5 reports the measured flux density or, in the absence of a detection, the upper limit on the flux density. Columns 6 and 7 give the sky coordinates (Right Ascension and Declination in $h$:$m$:$s$ and $d$:$m$:$s$, respectively) of the candidate EM counterpart, when provided. Column 8 shows the number of flags (see Section 7 for a detailed description); ‘NA’ marks entries for which GPT 4.1 never returned the ground truth even over multiple runs. A superscript asterisk () denotes a radio detection; all other entries represent non-detections (upper limits). The values in Columns 1–7 represent what would be produced by a trained expert through manual extraction and curation of GW170817-related GCN circulars.